Dissolving 120 g of urea (mol. wt. 60) in 1000 g of water gave a solution of density 1.15 g/mL. The


molarityof the solution is

plz giv a detaild xplanation!! 


Grade:11

4 Answers

zahoor ahmad
12 years ago
as density is given also molality can be calculated then we know m=molarity*1000/1000*density-molarity*molecular mass hence can be cal.
Hardik
6 years ago
We are given, DENSITY=1.5g/ml>MASS/VOLUME=1.15.............(1)NOW, MASS=1000+120=1120PUT THE VALUE IN EQN.1WE GET, VOLUME=973.91NOW, WE KNOWMOLARITY=NO.OF MOLES*1000/VOL.(IN ml)...(2)NO. OF MOLES = 120/60=2PUT IN EQN.2MOLARITY=2*1000/973.91=2.05.
varun
6 years ago
HELLO VIPUL GUPTA,
we know that,
density of the solution = mass of the solution/volume of the solution
therefore,volume of the solution = mass/density = 120 +1000/1.15 m/L
                                                                        =974X10^-3 litres
Now,,,
no of moles of solute = 120g/60.g 
                                 = 2 mol
(since mol wt of urea = 60)
Now,
Molarity = no of moles of solute/volume of solution in litres
Molarity = 2/974X10^-3
             =2000/974 
             = 2.05M
THANK YOU
Kali
6 years ago
Density of the solution 1.15gmL−1=MV1.15gmL−1=MV=1000+120V=1000+120V→V=11201.15→V=11201.15=973.91mL=973.91mL⇒⇒ Molarity =WM=WM×1000VmL×1000VmL=120×1000×1.2560×973.91=120×1000×1.2560×973.91=2.05M